Shift summary for owl, 12/17/99

Ed Blucher (blucher@hep.uchicago.edu)
Fri, 17 Dec 1999 06:59:34 -0600


Shift Summary for Owl Shift, 12/17/99

On Shift: Greg Bock, Ed Blucher

Hours of beam available at all intensities: 7/7
Hours of beam available at nominal intensity: 7
Hours of E799 data (special runs not included): 6.7
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time:
start/stop run, tape change, pipeline reset

========================================================================

Runs ended this shift:
15236 BEAM_799 (22.5 M events)
15237 BEAM_799 (in progress)

Tapes brought to FCC this shift:
UPA,B,C 229-231, UPS077

========================================================================
Overview of shift:
799 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
0:30 Scaler alarm limits relaxed slightly for BA1, SA4,
RC8, CIA.

15236 505 2:20 L3 accept fraction alarm. Stop run.
Install ADGN constants following run.

15237 2:40 Start run

6:50 Pipeline errors. Reset pipeline
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift